OpenStudy (debpriya):

The product of all the factors of a number N is equal to the number raised to the power of four. If N is less than 100 then find the possible values of N .( A. 5 B. 6 C. 9 D .10)

ganeshie8 (ganeshie8):

As a start, notice that if \(d\) is a factor of \(N\), then \(\dfrac{N}{d}\) is also a factor of \(N\).

ganeshie8 (ganeshie8):

As \(d\) runs over all the factors of \(N\), the expression \(\dfrac{N}{d}\) also runs over all the factor of \(N\)

ganeshie8 (ganeshie8):

Consider the product of all the factors \(d\) and \(N/d \) : \[ \left(\prod d \right)\left(\prod \dfrac{N}{d} \right) = \left(\prod d*\dfrac{N}{d} \right)= \left(\prod N \right) = N^{\tau(N)} \]

ganeshie8 (ganeshie8):

where \(\tau(N)\) is the number of factors of \(N\)

ganeshie8 (ganeshie8):

basically you need to solve the equation : \[N^4 = N^{\tau(N)/2}\]

ganeshie8 (ganeshie8):

that is same as solving \[\tau(N) = 8\]

ganeshie8 (ganeshie8):

how many positive integers less than 100 have exactly 8 factors ?

OpenStudy (debpriya):

Yes got it ...so after that how will we find the numbers less than 100 having exactly 8 factors ?

ganeshie8 (ganeshie8):

yea that looks tricky..

ganeshie8 (ganeshie8):

Consider the prime factorization of \(N\)

ganeshie8 (ganeshie8):

Would you agree that \(N\) cannot have more than \(3\) different prime factors

OpenStudy (debpriya):

Yes yes

ganeshie8 (ganeshie8):

so N can be of below forms : \(2^a3^b5^c\) such that \(a+b+c=8\)

OpenStudy (debpriya):

But why only 2,3 5, ? Why cant we take 5,7,11?

ganeshie8 (ganeshie8):

Oh right, we should consider all small primes

OpenStudy (debpriya):

Oh the product will be more than 100!

OpenStudy (debpriya):

Got it ! Thank you so much for your help !

ganeshie8 (ganeshie8):

may i see how you work the final answer

OpenStudy (debpriya):

8 can be written as 1+7 , 2+6, 3+5, 4+4 which can also be 1+6+1,1+2+5,1+3+4,1+1+5,1+4+2,3+4+1,3+2+3,2+1+5,2+2+4 ?

OpenStudy (debpriya):

Is there a shorter method ?

ganeshie8 (ganeshie8):

I made a mistake

ganeshie8 (ganeshie8):

\(N\) can be of form \(2^a3^b5^c7^d11^e\) such that \((a+1)(b+1)(c+1)(d+1)(e+1)=8\) right ?

ganeshie8 (ganeshie8):

because we want the number of factors to be equal to \(8\) not the exponents of primes

ganeshie8 (ganeshie8):

here are the possible N values [24,30,40,42,54,56,66,70,78,88]

OpenStudy (debpriya):

Apologies for late reply since internet connection is slow

OpenStudy (debpriya):

Why have we considered 5,7,11 because won't the product be greater than 100 and doesn't the question mention that N is lesser than 100 ?